OpenStudy (anonymous):

Dividing Radical Expressions Find the value of 7 times x squared all over the square root of x to the fifth power. please show steps and explain

OpenStudy (anonymous):

would that be 7 square root x

OpenStudy (campbell_st):

you can use the index laws for this problem 1st part \[\sqrt{x^5} = x^{\frac {5}{2}}\] \[\frac{1}{\sqrt{x^5}} = x^{-\frac{5}{2}}\] \[\frac{7x^2}{\sqrt{x^5}} = 7x^2 \times x^{-\frac{5}{2}}\] now using the multiplication law \[x^a \times x ^b = x^{a +b}\] \[7 \times x^2 \times x^{-\frac{5}{2}} = 7x^{2 -\frac{5}{2}} = 7x^{-\frac{1}{2}}\] which can be written \[\frac{7}{\sqrt{x}}\] by rationalising the denominator you get \[\frac{7\sqrt{x}}{x}\]

OpenStudy (anonymous):

thank you so much. i kinda have a stupid question but how would i find the exact value of the square root of 2 over the square root of 8. i know square root of 8 does into 4 x 2 but does 4 break down again o does it stay the same

OpenStudy (campbell_st):

\[\frac{\sqrt{2}}{2\sqrt{2}} = = \frac{1}{2} \times \frac{\sqrt{2}}{\sqrt{2}} = \frac{1}{2}\]

OpenStudy (anonymous):

thank you